Problem 2
One kg of air initially at 5 bar and 350 K and 3 kg of carbon dioxide (CO2), initially at 2 bar and 450 K, are located in opposite compartments within a well-insulated rigid tank, separated by a movable wall free. The wall also allows heat to pass through without energy being stored in the partition wall. Both air and CO2 behave like ideal gases. Determine the final equilibrium temperature, in K, and the final pressure, in bar, considering the specific heats constant.

What's missing?
Answer #1

air 5) 5 bar pressure , 350 k temperature, 2 kg W = 3kg, 2 bar pressure, 450K temperature : lequilibrium temperature is detervolume of air=? Air: p=5bar=580.987 atm = 4.935 atm м = 10 - 1000 M = 28.9647 = 29gmat (O2 and Ne major components) R=0.0821Jinsilarly - volume of cod=? Ne net - (M) x RXT p . = (3000 ) x 0°0821 atyremy x 24508 ( 4 4 gurex ? o-987 x 2 aty 127 6.07423-1 atm p = 102.665 x 0:08217 1476.861 P = 2.41473 atus p ~ 2.415 atm Ibar =0.987 atm - 1 atm = 1 bar 0.987 so 20415 atm =